OpenStudy (anonymous):

What is the domain of the function F(x) graphed below? f(x)=1/(x-1)

OpenStudy (anonymous):

A. x 1 B. -1 < x < 1 C. x 1 D. All real numbers

OpenStudy (anonymous):

PLEASE HELP

OpenStudy (lgbasallote):

\(x - 1 \ne 0\)

OpenStudy (lgbasallote):

therefore \(x \ne 1\)

OpenStudy (binary3i):

x not equal to 1 and all other numbers.

OpenStudy (callisto):

all real x except x=1

OpenStudy (callisto):

when x=1. x-1 = 1-1 =0 => the fraction is undefined, so, x=/=1 for the function

OpenStudy (anonymous):

Thanks so much you guys ! Big Help !!!!!!!!! All of you